# Group-Agent Reinforcement Learning with Heterogeneous Agents

### v1

@@ 1 Introduction
There has been an increasing interest in reinforcement learning involving multiple agents, emphasizing and strengthening different aspects of learning. One mainstream of such research is multi-agent reinforcement learning (MARL) Christianos et al. [2020], Sunehag et al. [2018], Mahajan et al. [2019], Wang et al. [2020], Rashid et al. [2018], Lowe et al. [2017]. It emphasizes cooperation and/or competition between multiple agents, in addition to individual success of each agent, in a shared environment with all the agents acting synchronously. The behaviour of cooperation or competition plays an important role, given that the outcome of each agent’s behaviour is affected by the behaviour of the other agents. MARL has been shown to improve the learning for playing the game of Go Silver et al. [2016, 2017, 2018], Schrittwieser et al. [2020] and robotic soccer, with the former being a competitive scenario and the later being a mixed task of cooperation and competition. Another mainstream is ensemble reinforcement learning (ERL) which aims at achieving better performance by combining the learned policies of multiple independent agents of either the same type Faußer and Schwenker [2011], Duell and Udluft [2013], Faußer and Schwenker [2015a, b], Osband et al. [2016], Saphal et al. [2021], Lee et al. [2021], Chen et al. [2017], Anschel et al. [2017], Smit et al. [2021] or different types Yang et al. [2020], Wiering and Van Hasselt [2008], Chen et al. [2018], Németh and Szűcs [2022], referred to as base agents. There is no communication between the base agents, while their decisions are combined only after completing the whole learning process. ERL does not intend to improve each individual base agent, but aims at a robust ensemble, of which an example application is stock trading Yang et al. [2020], Németh and Szűcs [2022].
@@
More recently, the concept of group-agent reinforcement learning (GARL) has been proposed by Wu and Zeng [2023]. It takes the inspiration from social learning where people learn by observing and imitating the behaviour of the others McLeod [2011]. In addition to learning by the classical trial-and-error exploration, the agents share knowledge with each other during the learning process. The goal is to enable the improvement of each individual agent, i.e., their learning speed and quality, through their received knowledge. Different from MARL, each GARL agent acts separately in its own environment, and learns asynchronously without any intention to cooperate or compete. This asynchronous group-learning paradigm is suitable for applications involving a group of agents with its members seeking to share and improve knowledge whenever available, e.g. video game playing, autonomous driving, network routing, etc.
@@
Given the different learning objectives and assumptions imposed by MARL, ERL and GARL, it is not possible to transfer effortlessly the technical success between them. We will review main advances in MARL and ERL, and discuss in comparison to GARL in more detail in Section 2. Significant research efforts have been put separately on improving each of these learning paradigms, addressing specific learning requirements. In this paper, we focus on advancing the GARL scenario, which is a new learning paradigm still under development. Currently, GARL Wu and Zeng [2023] has successfully improved the stability and scalability of agent training for a classic control task of CartPole-v0 on OpenAI Gym Brockman et al. [2016]. However, it works under a simple homogeneous setting, where the agents are restricted to act in the same environment and learn by the same algorithm. When being required to support agents that use different learning algorithms, such method becomes inapplicable, limited by its core design of communicating knowledge by sharing gradients.
@@
In this paper, we focus on developing more general GARL learning mechanisms to support heterogeneous agents underpinned by different algorithms, referred to as heterogeneous group-agent reinforcement learning (HGARL). We specify the knowledge to share between agents as their policy/value model parameters and their accumulated reward score. Each agent selects the best one from a set of actions suggested by not only their own policy model but also the received policy models from the other agents, for which three action selection rules are suggested, including probability addition (PA), probability multiplication (PM) and reward-value-likelihood combination (Combo). After this, the agent applies the selected action and continues its learning from the trajectory resulted from this selected action. In addition, we allow each agent to replace their own policy model with a better received model guided by a set of model adoption rules. Extensive experiments are conducted to assess the proposed approach, through learning to play a total of 43 different Atari 2600 games Bellemare et al. [2013]. We simulate a group of three agents each supported by a different Actor-Critic deep reinforcement learning algorithm, including A2C Mnih et al. [2016], PPO Schulman et al. [2017] and ACER Wang et al. [2017]. The effectiveness of HGARL is evidenced by performance improvement of these three types of agents, comparing learning in a group as opposed to learning individually. Among the total 129 agents examined, 96% are able to achieve a learning speed-up, and 72% learn over 100 times faster. Also, around 41% of the agents have achieved a higher reward score within less than 5% of the time steps required by a single agent when learning on its own. Overall, the contribution of this paper is three-fold:
@@ 2.1 Multi-Agent Reinforcement Learning
Example work from the first non-communication group include Sunehag et al. [2018] that decomposes the team value function into a sum of agent-wise value functions, Rashid et al. [2018] that relaxes the strong additive assumption of Sunehag et al. [2018], Mahajan et al. [2019] that improves Sunehag et al. [2018] as it severely limits the complexity of the team value function, and Wang et al. [2020] that applies the role concept to MARL to allow agents with similar roles to share similar behaviours. Another representative work from this group is Christianos et al. [2020], which shares experiences between multiple agents and incorporates all those into the loss function. All the agents are synchronised in a common environment, thus all of them are at the same learning stage at any time, and their decisions can always help each other. However, we have observed through experiments that such a synchronised sharing of experiences does not suit the GARL scenario. This is because our agents are normally in different learning stages and sharing the activities of the agents that are less mature to a senior agent will lead to potential learning failure.
For the second group of communicative approaches, all the agents are assumed to synchronise with each other when taking actions in a common environment Lowe et al. [2017], Sukhbaatar et al. [2016], Peng et al. [2017], Wang and Sartoretti [2022], Qu et al. [2019]. For instance, Lowe et al. [2017] considers explicit communication between agents but does not specify any particular structure on the communication method. Qu et al. [2019] explores a fully decentralised setting for the agents and locates each of them at a node of a communication network. Sukhbaatar et al. [2016] maintains a centralised communication network and allows multiple continuous communication cycles at each time step to decide the actions of all the agents. Peng et al. [2017] further deals with heterogeneous agents with bi-directional communication channel using recurrent neural networks. Wang and Sartoretti [2022] focuses on the problem of learning to communicate, introducing a new framework for the agents to learn a communication protocol.
@@ 2.2 Ensemble Reinforcement Learning
Ensemble learning is popular for its capacity of combining machine learning models to make a better prediction. It is well explored under the supervised learning paradigm, with well-known methods including bagging Breiman [1996], AdaBoost Freund et al. [1999] and random forest Ho [1995]. More recently, its effectiveness has been also demonstrated in reinforcement learning, through combining multiple policies learned by one common reinforcement learning algorithm Faußer and Schwenker [2011], Duell and Udluft [2013], Faußer and Schwenker [2015a, b], Osband et al. [2016], Saphal et al. [2021], Lee et al. [2021], Chen et al. [2017], Anschel et al. [2017], Smit et al. [2021], or combining multiple algorithms [Yang et al., 2020, Wiering and Van Hasselt, 2008, Chen et al., 2018, Németh and Szűcs, 2022].
To combine policies, the commonly used combiners include majority voting and averaging, e.g. for combining base value functions that are either parameterised or approximated by neural networks Faußer and Schwenker [2011, 2015a, 2015b], for combining policies in discrete action space Duell and Udluft [2013], and for developing an ensemble version of DQN algorithm Mnih et al. [2013] by averaging the Q-value functions Anschel et al. [2017]. There are also approaches developed to increase diversity between base policies, e.g., bootstrapping by random initialisation of policy networks Osband et al. [2016], exploration strategy based on confidence bounds Chen et al. [2017], and hybrid method for combining the previous two Lee et al. [2021], as well as the directed policy perturbation strategy Saphal et al. [2021] and the method for dealing with value overestimation Smit et al. [2021]. To combine algorithms, in addition to the standard approaches like majority voting, rank voting, Boltzmann multiplication and Boltzmann addition Wiering and Van Hasselt [2008], Chen et al. [2018] has developed a two-step ensemble strategy to combine DQN algorithms Mnih et al. [2013]. More specialized ensemble strategies have also been developed for stock trading Yang et al. [2020], Németh and Szűcs [2022].
@@ 3.1 Agent Simulation
We simulate different agents using different Actor-Critic reinforcement learning algorithms, selecting A2C Mnih et al. [2016], PPO Schulman et al. [2017] and ACER Wang et al. [2017] without loss of generality.
@@
Actor-Critic with Experience Replay (ACER) introduces several innovations into an off-policy Actor-Critic reinforcement learning algorithm. It adopts the Retrace algorithm Munos et al. [2016] in the estimation of the Q-function of the target policy. A Retrace estimator is expressed as
@@
where $\overline{\rho}_{t}=\min\{c,\rho_{t}\}$, $\rho_{t}=\frac{\pi(a_{t}|s_{t})}{\mu(a_{t}|s_{t})}$ with $\pi$ as the target policy while $\mu$ as the behaviour policy, $Q$ is the current estimate of the action value, and $V$ is the current estimate of the state value. $Q^{\textmd{ret}}$ is used in gradient calculation, as $(Q^{\textmd{ret}}(s_{t},a_{t})-Q(s_{t},a_{t}))\nabla Q(s_{t},a_{t})$. ACER adopts the importance weighted policy gradient that is approximated using marginal value functions over the limiting distribution of the learning process Degris et al. [2012]. It applies a truncation technique to the importance weights for the purposes of bounding the variance of the gradient estimate and correcting bias.
@@ 4.1 Experiment Setting
We test intensively the proposed HGARL on a total of 43 different Atari 2600 games Bellemare et al. [2013], and repeat the experiment 4 times for each game with 4 different seeds to obtain a more robust view on the learning performance. As explained in Section 3.1, we simulate a group of 3 agents supported by the learning algorithms of A2C, PPO and ACER, respectively. The goal is to test whether and how much HGARL improves the learning performance of each agent by letting them learn in a group, compared to learn on their own. We experimented with all the three selection rules, i.e., PA, PM and Combo, for HGARL. The experiments consume, for each run, 64 CPU cores with a total memory of 384GB for at least 7 days, which accounts for more than 10M core hours in total.
We used the A2C, PPO and ACER implementations in Dhariwal et al. [2017] released under MIT license. The ACER implementation learns an action value function, while the A2C and PPO implementations learn a state value function, where convolutional neural networks consisting of three convolutional layers with ReLU activation followed by one linear layer are used to model the policy and value functions. As a result, in the model adoption step, A2C and PPO agents can adopt both the policy and value networks (essentially their parameters) from each other, but ACER agent adopts only the policy network from A2C or PPO agent and vice versa. To implement the Combo rule, A2C and PPO predict the state values used in Eq. (7) by their value networks, while for the ACER agent, this value is computed from the predicted action probabilities by $V(s)=\sum_{a}\pi(a|s)\cdot Q(s,a)$, where $Q(s,a)$ is the action value, $V(s)$ is the state value and $\pi(a|s)$ is the policy. The batch size is set as 5 for A2C, 2048 for PPO and 20 for ACER, which is in general a sensitive parameter affecting the learning performance. These settings are kept as the original settings in Dhariwal et al. [2017] in order to get best tuned performance. The threshold $\phi$ for the negative log likelihood is set as $80\%$ of the initial negative log likelihood, which is fine-tuned.
@@ 5 Conclusion
GARL is a recently proposed concept which identifies an unexplored area in reinforcement learning. This work extends the method in Wu and Zeng [2023] which is restricted in a scenario where the agents are of the same type and in same environments to a more advanced approach which works with different types of agents in same environments. The evaluations show that we significantly improved the learning speed and quality of the agents in a GARL system compared to single-agent learning.

### v2

@@ 1 Introduction
There has been an increasing interest in reinforcement learning involving multiple agents, emphasizing and strengthening different aspects of learning. One mainstream of such research is multi-agent reinforcement learning (MARL) [Christianos et al., 2020, Sunehag et al., 2018, Mahajan et al., 2019, Wang et al., 2020, Rashid et al., 2018, Lowe et al., 2017]. It emphasizes cooperation and/or competition between multiple agents, in addition to individual success of each agent, in a shared environment with all the agents acting synchronously. The behaviour of cooperation or competition plays an important role, given that the outcome of each agent’s behaviour is affected by the behaviour of the other agents. MARL has been shown to improve the learning for playing the game of Go [Silver et al., 2016, 2017, 2018, Schrittwieser et al., 2020] and robotic soccer, with the former being a competitive scenario and the later being a mixed task of cooperation and competition. Another mainstream is ensemble reinforcement learning (ERL) which aims at achieving better performance by combining the learned policies of multiple independent agents of either the same type [Faußer and Schwenker, 2011, Duell and Udluft, 2013, Faußer and Schwenker, 2015a, b, Osband et al., 2016, Saphal et al., 2021, Lee et al., 2021, Chen et al., 2017, Anschel et al., 2017, Smit et al., 2021] or different types [Yang et al., 2020, Wiering and Van Hasselt, 2008, Chen et al., 2018, Németh and Szűcs, 2022], referred to as base agents. There is no communication between the base agents, while their decisions are combined only after completing the whole learning process. ERL does not intend to improve each individual base agent, but aims at a robust ensemble, of which an example application is stock trading [Yang et al., 2020, Németh and Szűcs, 2022].
@@
More recently, the concept of group-agent reinforcement learning (GARL) has been proposed by Wu and Zeng [2023]. It takes the inspiration from social learning where people learn by observing and imitating the behaviour of the others [McLeod, 2011]. In addition to learning by the classical trial-and-error exploration, the agents share knowledge with each other during the learning process. The goal is to enable the improvement of each individual agent, i.e., their learning speed and quality, through their received knowledge. Different from MARL, each GARL agent acts separately in its own environment, and learns asynchronously without any intention to cooperate or compete. This asynchronous group-learning paradigm is suitable for applications involving a group of agents with its members seeking to share and improve knowledge whenever available, e.g. video game playing, autonomous driving, network routing, etc.
@@
Given the different learning objectives and assumptions imposed by MARL, ERL and GARL, it is not possible to transfer effortlessly the technical success between them. We will review main advances in MARL and ERL, and discuss in comparison to GARL in more detail in Section 2. Significant research efforts have been put separately on improving each of these learning paradigms, addressing specific learning requirements. In this paper, we focus on advancing the GARL scenario, which is a new learning paradigm still under development. Currently, GARL [Wu and Zeng, 2023] has successfully improved the stability and scalability of agent training for a classic control task of CartPole-v0 on OpenAI Gym [Brockman et al., 2016]. However, it works under a simple homogeneous setting, where the agents are restricted to act in the same environment and learn by the same algorithm. When being required to support agents that use different learning algorithms, such method becomes inapplicable, limited by its core design of communicating knowledge by sharing gradients.
@@
In this paper, we focus on developing more general GARL learning mechanisms to support heterogeneous agents underpinned by different algorithms, referred to as heterogeneous group-agent reinforcement learning (HGARL). We specify the knowledge to share between agents as their policy/value model parameters and their accumulated reward score. Each agent selects the best one from a set of actions suggested by not only their own policy model but also the received policy models from the other agents, for which three action selection rules are suggested, including probability addition (PA), probability multiplication (PM) and reward-value-likelihood combination (Combo). After this, the agent applies the selected action and continues its learning from the trajectory resulted from this selected action. In addition, we allow each agent to replace their own policy model with a better received model guided by a set of model adoption rules. Extensive experiments are conducted to assess the proposed approach, through learning to play a total of 43 different Atari 2600 games [Bellemare et al., 2013]. We simulate a group of three agents each supported by a different Actor-Critic deep reinforcement learning algorithm, including A2C [Mnih et al., 2016], PPO [Schulman et al., 2017] and ACER [Wang et al., 2017]. The effectiveness of HGARL is evidenced by performance improvement of these three types of agents, comparing learning in a group as opposed to learning individually. Among the total 129 agents examined, 96% are able to achieve a learning speed-up, and 72% learn over 100 times faster. Also, around 41% of the agents have achieved a higher reward score within less than 5% of the time steps required by a single agent when learning on its own. Overall, the contribution of this paper is three-fold:
@@ 2.1 Multi-Agent Reinforcement Learning
Example work from the first non-communication group include [Sunehag et al., 2018] that decomposes the team value function into a sum of agent-wise value functions, [Rashid et al., 2018] that relaxes the strong additive assumption of [Sunehag et al., 2018], [Mahajan et al., 2019] that improves [Sunehag et al., 2018] as it severely limits the complexity of the team value function, and [Wang et al., 2020] that applies the role concept to MARL to allow agents with similar roles to share similar behaviours. Another representative work from this group is [Christianos et al., 2020], which shares experiences between multiple agents and incorporates all those into the loss function. All the agents are synchronised in a common environment, thus all of them are at the same learning stage at any time, and their decisions can always help each other. However, we have observed through experiments that such a synchronised sharing of experiences does not suit the GARL scenario. This is because our agents are normally in different learning stages and sharing the activities of the agents that are less mature to a senior agent will lead to potential learning failure.
For the second group of communicative approaches, all the agents are assumed to synchronise with each other when taking actions in a common environment [Lowe et al., 2017, Sukhbaatar et al., 2016, Peng et al., 2017, Wang and Sartoretti, 2022, Qu et al., 2019]. For instance, Lowe et al. [2017] consider explicit communication between agents but does not specify any particular structure on the communication method. Qu et al. [2019] explore a fully decentralised setting for the agents and locates each of them at a node of a communication network. Sukhbaatar et al. [2016] maintain a centralised communication network and allows multiple continuous communication cycles at each time step to decide the actions of all the agents. Peng et al. [2017] further deal with heterogeneous agents with bi-directional communication channel using recurrent neural networks. Wang and Sartoretti [2022] focus on the problem of learning to communicate, introducing a new framework for the agents to learn a communication protocol.
@@ 2.2 Ensemble Reinforcement Learning
Ensemble learning is popular for its capacity of combining machine learning models to make a better prediction. It is well explored under the supervised learning paradigm, with well-known methods including bagging [Breiman, 1996], AdaBoost [Freund et al., 1999] and random forest [Ho, 1995]. More recently, its effectiveness has been also demonstrated in reinforcement learning, through combining multiple policies learned by one common reinforcement learning algorithm [Faußer and Schwenker, 2011, Duell and Udluft, 2013, Faußer and Schwenker, 2015a, b, Osband et al., 2016, Saphal et al., 2021, Lee et al., 2021, Chen et al., 2017, Anschel et al., 2017, Smit et al., 2021], or combining multiple algorithms [Yang et al., 2020, Wiering and Van Hasselt, 2008, Chen et al., 2018, Németh and Szűcs, 2022].
To combine policies, the commonly used combiners include majority voting and averaging, e.g. for combining base value functions that are either parameterised or approximated by neural networks [Faußer and Schwenker, 2011, 2015a, 2015b], for combining policies in discrete action space [Duell and Udluft, 2013], and for developing an ensemble version of DQN algorithm [Mnih et al., 2013] by averaging the Q-value functions [Anschel et al., 2017]. There are also approaches developed to increase diversity between base policies, e.g., bootstrapping by random initialisation of policy networks [Osband et al., 2016], exploration strategy based on confidence bounds [Chen et al., 2017], and hybrid method for combining the previous two [Lee et al., 2021], as well as the directed policy perturbation strategy [Saphal et al., 2021] and the method for dealing with value overestimation [Smit et al., 2021]. To combine algorithms, in addition to the standard approaches like majority voting, rank voting, Boltzmann multiplication and Boltzmann addition [Wiering and Van Hasselt, 2008], Chen et al. [2018] have developed a two-step ensemble strategy to combine DQN algorithms [Mnih et al., 2013]. More specialized ensemble strategies have also been developed for stock trading [Yang et al., 2020, Németh and Szűcs, 2022].
@@ 3.1 Agent Simulation
We simulate different agents using different Actor-Critic reinforcement learning algorithms, selecting A2C [Mnih et al., 2016], PPO [Schulman et al., 2017] and ACER [Wang et al., 2017] without loss of generality.
@@
Actor-Critic with Experience Replay (ACER) introduces several innovations into an off-policy Actor-Critic reinforcement learning algorithm. It adopts the Retrace algorithm [Munos et al., 2016] in the estimation of the Q-function of the target policy. A Retrace estimator is expressed as
@@
where $\overline{\rho}_{t}=\min\{c,\rho_{t}\}$, $\rho_{t}=\frac{\pi(a_{t}|s_{t})}{\mu(a_{t}|s_{t})}$ with $\pi$ as the target policy while $\mu$ as the behaviour policy, $Q$ is the current estimate of the action value, and $V$ is the current estimate of the state value. $Q^{\textmd{ret}}$ is used in gradient calculation, as $(Q^{\textmd{ret}}(s_{t},a_{t})-Q(s_{t},a_{t}))\nabla Q(s_{t},a_{t})$. ACER adopts the importance weighted policy gradient that is approximated using marginal value functions over the limiting distribution of the learning process [Degris et al., 2012]. It applies a truncation technique to the importance weights for the purposes of bounding the variance of the gradient estimate and correcting bias.
@@ 4.1 Experiment Setting
We test intensively the proposed HGARL on a total of 43 different Atari 2600 games [Bellemare et al., 2013], and repeat the experiment 4 times for each game with 4 different seeds to obtain a more robust view on the learning performance. As explained in Section 3.1, we simulate a group of 3 agents supported by the learning algorithms of A2C, PPO and ACER, respectively. The goal is to test whether and how much HGARL improves the learning performance of each agent by letting them learn in a group, compared to learn on their own. We experimented with all the three selection rules, i.e., PA, PM and Combo, for HGARL. The experiments consume, for each run, 64 CPU cores with a total memory of 384GB for at least 7 days, which accounts for more than 10M core hours in total.
We used the A2C, PPO and ACER implementations by Dhariwal et al. [2017] released under MIT license. The ACER implementation learns an action value function, while the A2C and PPO implementations learn a state value function, where convolutional neural networks consisting of three convolutional layers with ReLU activation followed by one linear layer are used to model the policy and value functions. As a result, in the model adoption step, A2C and PPO agents can adopt both the policy and value networks (essentially their parameters) from each other, but ACER agent adopts only the policy network from A2C or PPO agent and vice versa. To implement the Combo rule, A2C and PPO predict the state values used in Eq. (7) by their value networks, while for the ACER agent, this value is computed from the predicted action probabilities by $V(s)=\sum_{a}\pi(a|s)\cdot Q(s,a)$, where $Q(s,a)$ is the action value, $V(s)$ is the state value and $\pi(a|s)$ is the policy. The batch size is set as 5 for A2C, 2048 for PPO and 20 for ACER, which is in general a sensitive parameter affecting the learning performance. These settings are kept as the original settings in [Dhariwal et al., 2017] in order to get best tuned performance. The threshold $\phi$ for the negative log likelihood is set as $80\%$ of the initial negative log likelihood, which is fine-tuned.
@@ 5 Conclusion
GARL is a new concept which identifies an unexplored area in reinforcement learning. This work extends the method in [Wu and Zeng, 2023] which is restricted in a scenario where the agents are of the same type and in same environments to a much more advanced approach which works in the more general learning scenario with different types of agents in same environments. The evaluations show that we significantly improved the learning speed and quality of the agents in a GARL system compared to single-agent learning.